The length of a retngle is 1 in. more than twice its width. If the perimeter of the rectangle is 26 in., find the width of the rectangel
A) 4 in.
B) 5 in.
C) 6 in.
D) 3 in.

Answer by Rekham(3)

The length of a retngle is 1 in. more than twice its width. If the perimeter of the rectangle is 26 in., find the width of the rectangel
Ans:
L=Lenght
W=Width
W=w (suppose)
.
. . L=2w+1
2(L+W)=perimeter
=>2(2w+1+w)=26
=>2(3w+1)=26
=>3w+1=13
=>3w=12
=>w=4
Hence Ans is A